OpenStudy (anonymous):

\frac{ 4x-2 }{ a } = 3

OpenStudy (anonymous):

(4x-2)/x=3

OpenStudy (anonymous):

i need to find x

OpenStudy (anonymous):

\[\frac{ 4x-2 }{ x } = 3\]

OpenStudy (anonymous):

What happens when you multiply both sides by x to get rid of the denominator?

OpenStudy (anonymous):

you tell me :))

OpenStudy (anonymous):

a=2

OpenStudy (anonymous):

Correct! Good job!